# Thermal effects on warm chromoinflation

### v1

@@ Abstract
I Introduction
Warm inflation (WI) Berera:1995wh ; Berera:1995ie ; Berera:1998px has recently seen successful implementations in models featuring pseudo-Nambu-Goldstone scalar fields Bastero-Gil:2016qru ; Bastero-Gil:2019gao ; Berghaus:2019whh ; Laine:2021ego ; DeRocco:2021rzv . These models often involve axion-like fields directly coupled to non-Abelian gauge fields. The dissipation of the axion-inflaton into gauge fields can naturally lead to a thermal radiation bath, even from initial vacuum conditions Laine:2021ego ; DeRocco:2021rzv . This thermal bath is a hallmark of the warm inflationary regime.
The explicit realization of consistent warm inflation dynamics in these models is a significant achievement. WI has emerged as a promising inflation model that aligns with effective field theory and potentially finds a ultraviolet (UV) completion in quantum gravity Das:2018rpg ; Motaharfar:2018zyb ; Goswami:2019ehb ; Berera:2019zdd ; Kamali:2019xnt ; Berera:2020iyn ; Das:2020xmh ; Brandenberger:2020oav ; Motaharfar:2021egj . Therefore, constructing explicit models of warm inflation and exploring their implications has become increasingly important.
In this work, we investigate an axion warm inflation (WI) model incorporating a Chern-Simons interaction between the axion inflaton field and a $SU(2)$ gauge field. It is well-established Maleknejad:2011jw ; Maleknejad:2016qjz ; Adshead:2012kp ; Kamali:2019ppi that, in addition to the background inflaton field, an isotropic solution for the $SU(2)$ gauge field is also permissible (for a review, see, e.g. Ref. Maleknejad:2012fw ). This background gauge field can substantially influence the evolutionary dynamics of the expanding system at both the background and perturbation levels.
Here, we demonstrate that the presence of a thermal bath naturally induces a plasma mass for the background gauge field. This thermal mass contribution to the gauge background field, similar to the thermal mass affecting the inflaton equation in warm inflation Berera:1998gx ; Yokoyama:1998ju , can potentially disrupt the slow-roll conditions for the gauge field, driving it towards a vanishing value. We investigate the impact of these thermal effects on the background dynamics and illustrate our results using a commonly employed axion potential.
This paper is organized as follows. In Sec. II, we briefly review the background dynamics of chromoinflation in the context of warm inflation. In Sec. III, we compute the leading order thermal contributions to the background dynamics. In Sec. IV, we illustrate our results by explicitly studying the dynamics in a case of axion-like potential. In Sec. V, we give our concluding remarks.
@@ I Introduction
II The Axion Warm Inflation Model
@@
$S$ $=$ $\int d^{4}xa^{3}(t)\left[\frac{1}{2}(\partial_{\mu}\phi)^{2}-V(\phi)-\frac{1}{4}F_{\mu\nu}^{c}F^{c\,\mu\nu}\right.$ (1)
@@
$\tilde{F}^{c\,\mu\nu}=\frac{\epsilon^{\mu\nu\rho\sigma}}{2a^{3}(t)}F_{\rho\sigma}^{c},$ (2)
@@
The $SU(2)$ gauge field admit a homogeneous vacuum expectation value (VEV), defined as Maleknejad:2011jw ; Maleknejad:2016qjz ; Adshead:2012kp
$\bar{A}^{c}_{i}=a(t)\psi(t)\delta^{c}_{i},\;\;\;\;\bar{A}^{c}_{0}=0,$ (3)
@@
$\bar{F}_{0i}^{c}=-\bar{F}_{i0}^{c}=a(H\psi+\dot{\psi})\delta_{i}^{c},\;\;\;\bar{F}_{ij}^{c}=ga^{2}\psi^{2}\epsilon_{ij}^{c}.$ (4)
Note that in chromoinflation Maleknejad:2011jw ; Maleknejad:2016qjz ; Adshead:2012kp the coupling $\lambda$ between the inflaton $\phi$ and the gauge fields in the interaction term in Eq. (1) is assumed to be an independent constant. Typically, a consistent background evolution for the inflaton and gauge field is achieved for large couplings, or more generally, $\lambda M_{\rm Pl}/f\gg 1$. How to achieve such large couplings in effective theories has been discussed recently for example in Refs. Agrawal:2018mkd ; Holland:2020jdh . Note that in the axion dynamics case, the coupling $\lambda$ is simply related to the gauge coupling $g$ by $\lambda=g^{2}/(8\pi^{2})$.
The inflaton (axion) field has been shown to dissipate into radiation bath gauge fields, which forms a thermal bath with temperature $T$ and whose dissipation coefficient is given by Berghaus:2019whh ; Laine:2021ego
$\Upsilon(T)=\kappa\frac{T^{3}}{f^{2}},$ (5)
@@
$\kappa\simeq 1.2\pi\frac{g^{4}(g^{2}N_{c})^{3}(N_{c}^{2}-1)}{(64\pi^{3})^{2}}\left[\ln\left(\frac{m_{D}}{\gamma}\right)+3.041\right],$ (6)
where $m_{D}^{2}=g^{2}N_{c}T^{2}/3$ is the Debye mass squared of the Yang-Mills plasma and $\gamma$ is given by the solution of Moore:2010jd
$\gamma=\frac{g^{2}N_{c}T}{4\pi}\left[\ln\left(\frac{m_{D}}{\gamma}\right)+3.041\right].$ (7)
@@
$\kappa$ $=$ $0.3\alpha^{5}N_{c}^{3}(N_{c}^{2}-1)W\left(e^{3.041}\sqrt{\frac{4\pi}{3\alpha N_{c}}}\right),$ (8)
where $\alpha=g^{2}/(4\pi)$ is the fine structure Yang-Mills coupling and $W(x)$ is the Lambert function, given by the principal solution of $x=we^{w}$. For $N_{c}=2$ and assuming e.g. $\alpha=0.1$, we obtain that $\kappa\simeq 2.4\times 10^{-4}$.
II.1 Background Evolution
Implementations of warm inflation in the context of chromoinflation have been considered in some recent works Yeasmin:2022ncm ; Mukuno:2024yoa . At the background level, the energy density related to the inflaton field $\phi$, the VEV $\psi$ of the Yang-Mills gauge field and the thermal bath are, respectively, given by
$\rho_{\phi}=\frac{1}{2}\dot{\phi}^{2}+V(\phi),$ (9)
$\rho_{YM}=\frac{3}{2}\left(\dot{\psi}+H\psi\right)^{2}+\frac{3}{2}g^{2}\psi^{4},$ (10)
$\rho_{r}=C_{r}T^{4},$ (11)
@@ II.1 Background Evolution
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\ddot{\phi}+3H\dot{\phi}+V_{,\phi}+3\frac{\lambda}{f}g\psi^{2}(\dot{\psi}+H\psi)+\Upsilon\dot{\phi}=0,$ (12)
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\ddot{\psi}+3H\dot{\psi}+(2H^{2}+\dot{H})\psi+2g^{2}\psi^{3}-\frac{\lambda g}{f}\psi^{2}\dot{\phi}=0,$ (13)
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\dot{\rho}_{r}+4H\rho_{r}=\Upsilon\dot{\phi}^{2},$ (14)
where the dissipation coefficient $\Upsilon$ is given by Eq. (5).
The slow-roll variables can be defined as Maleknejad:2016qjz
$\epsilon_{\psi}=\frac{\dot{\psi}}{H\psi}~{}~{}~{}~{}~{}~{}\eta_{\psi}=-\frac{\ddot{\psi}}{H\psi}$ (15)
@@
In the slow-roll approximation, $\epsilon_{\psi}<1$, $\eta_{\psi}<1$, $\epsilon_{H}<1$ and out of the instability condition Adshead:2013nka ; Namba:2013kia ; Dimastrogiovanni:2012ew (where the gauge perturbations become tachyonic), i.e., $g^{2}\psi^{2}/(2H^{2})>1$, the evolution of $\psi$ in Eq. (13) is simplified to
$g\psi^{3}\simeq\frac{\lambda}{2f}\psi^{2}\dot{\phi}.$ (16)
@@
There are several consistency conditions that need to be satisfied such that the above equations can work properly in a warm inflation context. First, a thermal bath of particles need to be produced during inflation. This has been explicitly verified in Refs. Laine:2021ego ; DeRocco:2021rzv . Second, backreaction from the nonlinearities which are inherent of non-Abelian gauge fields need to be sufficient suppressed such that the effective theory leading to the evolution equations (12), (13) and (14) remain valid. This typically requires that we work with temperatures that remain smaller than the axion decay constant DeRocco:2021rzv , i.e., $T<f$. The warm inflation regime itself also requires $T>H$. So, we must ensure the hierarchy of energy scales: $H<T<f$. Third, to avoid the instabilities of the non-Abelian gauge field, we are required to have $g|\psi|/H>\sqrt{2}$. Forth, perturbativity also requires that the gauge coupling to be small, $\alpha<1$, such that we can also trust in the computation leading to the dissipation coefficient Eq. (5). Fifth, to ensure that there is a thermal bath of gauge particles, the temperature must remain above the confinement scale of the model Berghaus:2019whh . This implies in particular in the choice for the inflaton potential $V(\phi)$. The usual low energy infrared (IR) potential assumed for the axion cannot be used here and some ultraviolet potential need to be used. Some different potentials have been considered before, like an hybrid type of potential as used in Ref. Berghaus:2019whh . An exponential potential was considered in Ref. Das:2019acf . A simple quadratic potential for the inflaton was used in Ref. Laine:2021ego , while in Ref. DeRocco:2021rzv some axion monodromy type of potentials have been considered.
III Thermal contributions to the background dynamics
@@
To take into account the effects of the thermal bath on the background dynamics, we will work analogously to the loop expansion in quantum field theory. This starts by expanding the fields around their vacuum background expectation values. In this work in special, we will be looking how the fluctuations around the gauge field background $\psi$ will affect the dynamics. Hence, the gauge field in (1) is expanded like $A_{\mu}\to A^{\prime}_{\mu}=\bar{A}_{\mu}(t)+A_{\mu}$, where $\langle A^{\prime}_{\mu}\rangle=\bar{A}_{\mu}(t)$, with $\bar{A}_{\mu}(t)$ given by Eq. (3) and $\langle A_{\mu}\rangle=0$. Working up to one-loop order in the fluctuations of the gauge field (i.e., second order in the gauge field fluctuations), we have for instance that
$F^{a}_{\mu\nu}=\bar{F}_{\mu\nu}^{a}+\left(\bar{D}_{\mu}A_{\nu}\right)^{a}-\left(\bar{D}_{\nu}A_{\mu}\right)^{a}+g\epsilon^{abc}A_{\mu}^{b}A_{\nu}^{c},$ (17)
where $\bar{F}_{\mu\nu}^{a}$ is given by Eq. (4) and
$\bar{D}_{\mu}=\partial_{\mu}-igJ^{a}\bar{A}_{\mu}^{a},$ (18)
@@
is the covariant derivative expressed in terms of the background gauge field, with $J^{a}$ the generators of the $SU(2)$ group. Working in the adjunct representation of the $SU(2)$ gauge group, $(J_{a})_{bc}=\epsilon_{abc}$.
@@
Using Eq. (17) in (1), the action, up to second order in the gauge field fluctuations, then takes the form
$S\simeq S_{0}+\delta^{2}S,$ (19)
@@
$S_{0}$ $=$ $\int d^{4}xa^{3}(t)\left[\frac{1}{2}(\partial_{\mu}\phi)^{2}-V(\phi)-\frac{1}{4}\bar{F}_{\mu\nu}^{c}\bar{F}^{c\,\mu\nu}\right.$ (20)
@@
is the zeroth order action in the fluctuations and $\delta^{2}S$ is the action when expanding the fields up to the second-order in the fluctuations (see also Ref. Adshead:2013nka for a similar approach for deriving the perturbation equations in chromoinflation). The Yang-Mills and Chern-Simons contributions can then be expressed as
@@
We still need to choose a gauge fixing term. Here, it becomes convenient to fix a gauge that depends explicitly on the background gauge field, such that (see, e.g. Srednicki:2007qs )
$\delta^{2}S_{\rm gf}=\int d^{4}x\,a^{3}\left[-\frac{1}{2\alpha}(\bar{D}^{\mu}A^{\mu})^{a}(\bar{D}_{\nu}A_{\nu})^{a}\right],$ (22)
@@
$\delta^{2}S_{\rm ghost}=\int d^{4}x\,a^{3}\left[(\bar{D}^{\mu}\bar{\eta})^{a}(\bar{D}_{\mu}\eta)^{a}\right],$ (23)
@@
By substituting in Eq. (LABEL:S2YM2) the equations (3) and (4), we have many terms depending on the background field $\psi$. Many of them vanishes when performing the functional integration over the fluctuations (e.g. the Chern-Simons dependent terms which are total derivatives). For the terms effectively contributing, one notices that in the presence of a nonvanishing background gauge field, both the gauge field and the ghosts acquire a mass square, $m_{A}^{2}=m_{\eta}^{2}=2g^{2}\psi^{2}$. Their functional integration lead to an effective potential for $\psi$ which can be expressed as [^2] Note that curvature effects are expected to be negligible for masses larger than the Hubble scale, e.g. for $2g^{2}\psi^{2}>H^{2}$ and, furthermore, at finite temperature, the relevant momentum modes for the gauge fluctuations contributing for the loop integrals are hard momenta, $k\sim T$. In particular, this later condition is well satisfied in warm inflation by definition, where $T>H$. This justifies a Minskowski like derivation of the loop terms used to derive the result given by Eq. (25).
$\!\!\!\!\!\Delta V_{\rm eff}(\psi)$ $=$ $(N_{c}^{2}-1)\int\frac{d^{4}k_{E}}{(2\pi)^{4}}\ln\left(k_{E}^{2}+2g^{2}\psi^{2}\right),$ (25)
where $k_{E}$ refers to the (physical) momentum expressed in Euclidean spacetime. Diagrammatically, the terms contributing to the effective potential at leading order are displayed in Fig. 1.
@@
From the usual rules of finite temperature quantum field theory Kapusta:2006pm , $k_{E}^{2}=k_{4}^{2}+{\bf k}^{2}$ and $k_{4}\equiv\omega_{n}=2\pi nT$, $n\in\mathbb{Z}$ and $\omega_{n}$ are the Matsubara’s frequencies for bosons. The integrals over momenta are expressed as
$\int\frac{d^{4}k_{E}}{(2\pi)^{4}}\to T\sum_{k_{4}=\omega_{n}}\int\frac{d^{3}k}{\left(2\pi\right)^{3}}.$ (26)
At zero temperature, the momentum integral in Eq. (25) is divergent. The divergent terms can be handled as usual in perturbation theory in quantum field theory by adding the appropriate renormalization counterterms in the tree-level action Eq. (20). The relevant contributions for us are the finite temperature ones, which, after performing the Matsubara’s frequency sum in Eq. (25), we are left with the thermal contribution,
$\Delta V_{\rm eff}(\psi,T)$ $=$ $2(N_{c}^{2}-1)\frac{T^{4}}{2\pi^{2}}J_{B}(y),$ (27)
where Kapusta:2006pm
@@
with $y^{2}=2g^{2}\psi^{2}/T^{2}$ and the last line in Eq. (28) follows by using an expansion for $y\ll 1$. We could as well consider the additional terms in the expansion in Eq. (28). However, for our purposes of demonstrating the effects of the thermal contributions, the truncation of the series up to quadratic order for the thermal integral $J_{B}(y)$ suffices for us [^3] We have nevertheless explicitly verified in our numerical analysis that the higher order terms remain subdominant compared to the quadratic term in Eq. (28).. We then obtain for Eq. (25) the result
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\ddot{\phi}+3H\dot{\phi}+V_{,\phi}+3\frac{\lambda}{f}g\psi^{2}(\dot{\psi}+H\psi)+\Upsilon\dot{\phi}=0,$ (33)
@@
$+m_{\psi}^{2}(T)\psi-\frac{\lambda g}{f}\psi^{2}\dot{\phi}=0,$ (34)
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!T\dot{s}+3HTs=\Upsilon\dot{\phi}^{2},$ (35)
@@
$H^{2}$ $=$ $\frac{1}{3M_{\rm Pl}^{2}}\left[\frac{\dot{\phi}^{2}}{2}+V(\phi)+\frac{3}{2}(\dot{\psi}+H\psi)^{2}\right.$ (36)
@@
We can qualitatively already investigate the effect of the thermal mass term in the background equation for $\psi$. In the slow-roll approximation, from Eq. (34), we obtain that
$3H\dot{\psi}\simeq-\left[2H^{2}+m_{\psi}^{2}(T)\right]\psi-2g^{2}\psi^{3}+\frac{\lambda g}{f}\psi^{2}\dot{\phi}.$ (37)
If we associate the right-hand side in Eq. (37) as an analogous of the field derivative of an effective potential for $\psi$ (see, e.g. Ref. Fujita:2021eue for a similar approach in cold chromoinflation), then we can define the effective potential (and assuming $\dot{\phi}$ as constant),
$V_{\rm eff}(\psi,T)\simeq\left[2H^{2}+m_{\psi}^{2}(T)\right]\frac{\psi^{2}}{2}-\frac{\lambda g\dot{\phi}}{f}\frac{\psi^{3}}{3}+g^{2}\frac{\psi^{4}}{2}.$ (38)

### v2

@@ Abstract
1 Introduction
Warm inflation [1, 2, 3] has recently seen successful implementations in models featuring pseudo-Nambu-Goldstone scalar fields [4, 5, 6, 7, 8]. These models often involve axion-like fields directly coupled to non-Abelian gauge fields. The dissipation of the axion-inflaton into gauge fields can naturally lead to a thermal radiation bath, even from initial vacuum conditions [9, 7, 8]. This thermal bath is a hallmark of the warm inflationary regime.
The explicit realization of consistent warm inflation dynamics in these models is a significant achievement. Warm inflation has emerged as a promising inflation model that aligns with effective field theory and potentially finds a ultraviolet (UV) completion in quantum gravity [10, 11, 12, 13, 14, 15, 16, 17, 18]. Therefore, constructing explicit models of warm inflation and exploring their implications has become increasingly important.
In this work, we investigate an axion warm inflation model incorporating a Chern-Simons interaction between the axion inflaton field and a $SU(2)$ gauge field. It is well-established [19, 20, 21, 22] that, in addition to the background inflaton field, an isotropic solution for the $SU(2)$ gauge field is also permissible (for a review, see, e.g. ref. [23]). This background gauge field can substantially influence the evolutionary dynamics of the expanding system at both the background and perturbation levels.
Here, we demonstrate that the presence of a thermal bath naturally induces a plasma mass for the background gauge field. This thermal mass contribution to the gauge background field, similar to the thermal mass affecting the inflaton equation in warm inflation [24, 25], can potentially disrupt the slow-roll conditions for the gauge field, driving it towards a vanishing value. We investigate the impact of these thermal effects on the background dynamics and illustrate our results using a commonly employed axion potential.
This paper is organized as follows. In section 2, we briefly review the background dynamics of chromoinflation in the context of warm inflation. In section 3, we compute the leading order thermal contributions to the background dynamics. In section 4, we illustrate our results by explicitly studying the dynamics in a case of axion-like potential. In section 5, we give our concluding remarks.
@@ 1 Introduction
2 The Axion Warm Inflation Model
@@
$S$ $=$ $\int d^{4}xa^{3}(t)\left[\frac{1}{2}(\partial_{\mu}\phi)^{2}-V(\phi)-\frac{1}{4}F_{\mu\nu}^{c}F^{c\,\mu\nu}\right.$ (2.1)
@@
$\tilde{F}^{c\,\mu\nu}=\frac{\epsilon^{\mu\nu\rho\sigma}}{2a^{3}(t)}F_{\rho\sigma}^{c},$ (2.2)
@@
The $SU(2)$ gauge field admits a homogeneous vacuum expectation value (VEV), defined as [19, 20, 21]
$\bar{A}^{c}_{i}=a(t)\psi(t)\delta^{c}_{i},\;\;\;\;\bar{A}^{c}_{0}=0,$ (2.3)
@@
$\bar{F}_{0i}^{c}=-\bar{F}_{i0}^{c}=a(H\psi+\dot{\psi})\delta_{i}^{c},\;\;\;\bar{F}_{ij}^{c}=ga^{2}\psi^{2}\epsilon_{ij}^{c}.$ (2.4)
Note that in chromoinflation [19, 20, 21] the coupling $\lambda$ between the inflaton $\phi$ and the gauge fields in the interaction term in eq. (2.1) is assumed to be an independent constant. Typically, a consistent background evolution for the inflaton and gauge field is achieved for large couplings, or more generally, $\lambda M_{\rm Pl}/f\gg 1$. How to achieve such large couplings in effective theories has been discussed recently for example in refs. [26, 27]. Note that in the axion dynamics case, the coupling $\lambda$ is simply related to the gauge coupling $g$ by $\lambda=g^{2}/(8\pi^{2})$.
The inflaton (axion) field has been shown to dissipate into radiation bath gauge fields, which forms a thermal bath with temperature $T$ and whose dissipation coefficient is given by [6, 7]
$\Upsilon(T)=\kappa\frac{T^{3}}{f^{2}},$ (2.5)
@@
$\kappa\simeq 1.2\pi\frac{g^{4}(g^{2}N_{c})^{3}(N_{c}^{2}-1)}{(64\pi^{3})^{2}}\left[\ln\left(\frac{m_{D}}{\gamma}\right)+3.041\right],$ (2.6)
where $m_{D}^{2}=g^{2}N_{c}T^{2}/3$ is the Debye mass squared of the Yang-Mills plasma and $\gamma$ is given by the solution of [28]
$\gamma=\frac{g^{2}N_{c}T}{4\pi}\left[\ln\left(\frac{m_{D}}{\gamma}\right)+3.041\right].$ (2.7)
@@
$\kappa$ $=$ $0.3\alpha_{g}^{5}N_{c}^{3}(N_{c}^{2}-1)W\left(e^{3.041}\sqrt{\frac{4\pi}{3\alpha_{g}N_{c}}}\right),$ (2.8)
where $\alpha_{g}=g^{2}/(4\pi)$ is the fine structure Yang-Mills coupling and $W(x)$ is the Lambert function, given by the principal solution of $x=we^{w}$. For $N_{c}=2$ and assuming e.g. $\alpha_{g}=0.1$, we obtain that $\kappa\simeq 2.4\times 10^{-4}$.
2.1 Background Evolution
Implementations of warm inflation in the context of chromoinflation have been considered in some recent works [29, 30]. At the background level, the energy density related to the inflaton field $\phi$, the VEV $\psi$ of the Yang-Mills gauge field and the thermal bath are, respectively, given by
$\rho_{\phi}=\frac{1}{2}\dot{\phi}^{2}+V(\phi),$ (2.9)
$\rho_{YM}=\frac{3}{2}\left(\dot{\psi}+H\psi\right)^{2}+\frac{3}{2}g^{2}\psi^{4},$ (2.10)
$\rho_{r}=C_{r}T^{4},$ (2.11)
@@ 2.1 Background Evolution
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\ddot{\phi}+3H\dot{\phi}+V_{,\phi}+3\frac{\lambda}{f}g\psi^{2}(\dot{\psi}+H\psi)+\Upsilon\dot{\phi}=0,$ (2.12)
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\ddot{\psi}+3H\dot{\psi}+(2H^{2}+\dot{H})\psi+2g^{2}\psi^{3}-\frac{\lambda g}{f}\psi^{2}\dot{\phi}=0,$ (2.13)
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\dot{\rho}_{r}+4H\rho_{r}=\Upsilon\dot{\phi}^{2},$ (2.14)
where the dissipation coefficient $\Upsilon$ is given by eq. (2.5).
The slow-roll variables can be defined as [20]
$\epsilon_{\psi}=\frac{\dot{\psi}}{H\psi}~{}~{}~{}~{}~{}~{}\eta_{\psi}=-\frac{\ddot{\psi}}{H\psi}$ (2.15)
@@
In the slow-roll approximation, $\epsilon_{\psi}<1$, $\eta_{\psi}<1$, $\epsilon_{H}<1$ and out of the instability condition [31, 32, 33] (where the gauge perturbations become tachyonic), i.e., $g^{2}\psi^{2}/(2H^{2})>1$, the evolution of $\psi$ in eq. (2.13) is simplified to
$g\psi^{3}\simeq\frac{\lambda}{2f}\psi^{2}\dot{\phi}.$ (2.16)
@@
There are several consistency conditions that need to be satisfied such that the above equations can work properly in a warm inflation context. First, a thermal bath of particles needs to be produced during inflation. This has been explicitly verified in refs. [7, 8]. Second, backreaction from the nonlinearities which are inherent of non-Abelian gauge fields needs to be sufficient suppressed such that the effective theory leading to the evolution equations (2.12), (2.13) and (2.14) remains valid. This typically requires that we work with temperatures that remain smaller than the axion decay constant [8], i.e., $T<f$. The warm inflation regime itself also requires $T>H$. So, we must ensure the hierarchy of energy scales: $H<T<f$. Third, to avoid the instabilities of the non-Abelian gauge field, we are required to have $g|\psi|/H>\sqrt{2}$. Forth, perturbativity also requires that the gauge coupling to be small, $\alpha_{g}<1$, such that we can also trust in the computation leading to the dissipation coefficient eq. (2.5). Fifth, to ensure that there is a thermal bath of gauge particles, the temperature must remain above the confinement scale of the model [6]. This implies in particular in the choice for the inflaton potential $V(\phi)$. The usual low energy infrared (IR) potential assumed for the axion cannot be used here and some ultraviolet potential needs to be used. Some different potentials have been considered before, like an hybrid type of potential as used in ref. [6]. An exponential potential was considered in ref. [12]. A simple quadratic potential for the inflaton was used in ref. [7], while in ref. [8] some axion monodromy type of potentials have been considered.
3 Thermal contributions to the background dynamics
@@
To take into account the effects of the thermal bath on the background dynamics, we will work analogously to the loop expansion in quantum field theory. This starts by expanding the fields around their vacuum background expectation values. In this work in particular, we will be looking how the fluctuations around the gauge field background $\psi$ will affect the dynamics. Hence, the gauge field in (2.1) is expanded like $A_{\mu}\to A^{\prime}_{\mu}=\bar{A}_{\mu}(t)+A_{\mu}$, where $\langle A^{\prime}_{\mu}\rangle=\bar{A}_{\mu}(t)$, with $\bar{A}_{\mu}(t)$ given by eq. (2.3) and $\langle A_{\mu}\rangle=0$. Hence, we have for instance that
$F^{a}_{\mu\nu}=\bar{F}_{\mu\nu}^{a}+\left(\bar{D}_{\mu}A_{\nu}\right)^{a}-\left(\bar{D}_{\nu}A_{\mu}\right)^{a}+g\epsilon^{abc}A_{\mu}^{b}A_{\nu}^{c},$ (3.1)
where $\bar{F}_{\mu\nu}^{a}$ is given by eq. (2.4) and
$\bar{D}_{\mu}=\partial_{\mu}-igJ^{a}\bar{A}_{\mu}^{a},$ (3.2)
@@
is the covariant derivative expressed in terms of the background gauge field, with $J^{a}$ the generators of the $SU(2)$ group. Working in the adjoint representation of the $SU(2)$ gauge group, $(J_{a})_{bc}=\epsilon_{abc}$.
@@
Using eq. (3.1) in (2.1), the action, up to second order in the gauge field fluctuations, then takes the form
$S\simeq S_{0}+\delta^{2}S,$ (3.3)
@@
$S_{0}$ $=$ $\int d^{4}xa^{3}(t)\left[\frac{1}{2}(\partial_{\mu}\phi)^{2}-V(\phi)-\frac{1}{4}\bar{F}_{\mu\nu}^{c}\bar{F}^{c\,\mu\nu}\right.$ (3.4)
@@
is the zeroth order action in the fluctuations and $\delta^{2}S$ is the action when expanding the fields up to the second-order in the fluctuations (see also ref. [31] for a similar approach for deriving the perturbation equations in chromoinflation). The Yang-Mills and Chern-Simons contributions can then be expressed as
@@
We still need to choose a gauge fixing term. Here, it becomes convenient to fix a gauge that depends explicitly on the background gauge field, such that (see, e.g. [34])
$\delta^{2}S_{\rm gf}=\int d^{4}x\,a^{3}\left[-\frac{1}{2\alpha}(\bar{D}^{\mu}A^{\mu})^{a}(\bar{D}_{\nu}A_{\nu})^{a}\right],$ (3.6)
@@
$\delta^{2}S_{\rm ghost}=\int d^{4}x\,a^{3}\left[(\bar{D}^{\mu}\bar{\eta})^{a}(\bar{D}_{\mu}\eta)^{a}\right],$ (3.7)
@@
By substituting in eq. (LABEL:S2YM2) the equations (2.3) and (2.4), we have many terms depending on the background field $\psi$. Many of them vanish when performing the functional integration over the fluctuations (e.g. the Chern-Simons dependent terms which are total derivatives). For the terms effectively contributing, one notices that in the presence of a nonvanishing background gauge field, both the gauge field and the ghosts acquire a mass square, $m_{A}^{2}=m_{\eta}^{2}=2g^{2}\psi^{2}$. Their functional integration lead to an effective potential for $\psi$ which can be expressed as [^3] Note that curvature effects are expected to be negligible for masses larger than the Hubble scale, e.g. for $2g^{2}\psi^{2}>H^{2}$ and, furthermore, at finite temperature, the relevant momentum modes for the gauge fluctuations contributing for the loop integrals are hard momenta, $k\sim T$. In particular, this later condition is well satisfied in warm inflation by definition, where $T>H$. This justifies a Minskowski like derivation of the loop terms used to derive the result given by eq. (3.9).
$\!\!\!\!\!\Delta V_{\rm eff}(\psi)$ $=$ $(N_{c}^{2}-1)\int\frac{d^{4}k_{E}}{(2\pi)^{4}}\ln\left(k_{E}^{2}+2g^{2}\psi^{2}\right),$ (3.9)
where $k_{E}$ refers to the (physical) momentum expressed in Euclidean spacetime. Diagrammatically, the terms contributing to the effective potential at leading order are displayed in figure 1.
@@
From the usual rules of finite temperature quantum field theory [35], $k_{E}^{2}=k_{4}^{2}+{\bf k}^{2}$ and $k_{4}\equiv\omega_{n}=2\pi nT$, $n\in\mathbb{Z}$ and $\omega_{n}$ are the Matsubara’s frequencies for bosons. The integrals over momenta are expressed as
$\int\frac{d^{4}k_{E}}{(2\pi)^{4}}\to T\sum_{k_{4}=\omega_{n}}\int\frac{d^{3}k}{\left(2\pi\right)^{3}}.$ (3.10)
At zero temperature, the momentum integral in eq. (3.9) is divergent. The divergent terms can be handled as usual in perturbation theory in quantum field theory by adding the appropriate renormalization counterterms in the tree-level action eq. (3.4). The relevant contributions for us are the finite temperature ones, which, after performing the Matsubara’s frequency sum in eq. (3.9), we are left with the thermal contribution,
$\Delta V_{\rm eff}(\psi,T)$ $=$ $2(N_{c}^{2}-1)\frac{T^{4}}{2\pi^{2}}J_{B}(y),$ (3.11)
where [35]
@@
with $y^{2}=2g^{2}\psi^{2}/T^{2}$ and the last line in eq. (3.12) follows by using an expansion for $y\ll 1$. We could as well consider the additional terms in the expansion in eq. (3.12). However, for our purposes of demonstrating the effects of the thermal contributions, the truncation of the series up to quadratic order for the thermal integral $J_{B}(y)$ suffices for us [^4] We have nevertheless explicitly verified in our numerical analysis that the higher order terms remain subdominant compared to the quadratic term in eq. (3.12).. We then obtain for eq. (3.9) the result
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\ddot{\phi}+3H\dot{\phi}+V_{,\phi}+3\frac{\lambda}{f}g\psi^{2}(\dot{\psi}+H\psi)+\Upsilon\dot{\phi}=0,$ (3.17)
@@
$+m_{\psi}^{2}(T)\psi-\frac{\lambda g}{f}\psi^{2}\dot{\phi}=0,$ (3.18)
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!T\dot{s}+3HTs=\Upsilon\dot{\phi}^{2},$ (3.19)
@@
$H^{2}$ $=$ $\frac{1}{3M_{\rm Pl}^{2}}\left[\frac{\dot{\phi}^{2}}{2}+V(\phi)+\frac{3}{2}(\dot{\psi}+H\psi)^{2}\right.$ (3.20)
@@
We can qualitatively already investigate the effect of the thermal mass term in the background equation for $\psi$. In the slow-roll approximation, from eq. (3.18), we obtain that
$3H\dot{\psi}\simeq-\left[2H^{2}+m_{\psi}^{2}(T)\right]\psi-2g^{2}\psi^{3}+\frac{\lambda g}{f}\psi^{2}\dot{\phi}.$ (3.21)
If we associate the right-hand side in eq. (3.21) with an analogous of the field derivative of an effective potential for $\psi$ (see, e.g. ref. [36] for a similar approach in cold chromoinflation), then we can define the effective potential (and assuming $\dot{\phi}$ as constant),
$V_{\rm eff}(\psi,T)\simeq\left[2H^{2}+m_{\psi}^{2}(T)\right]\frac{\psi^{2}}{2}-\frac{\lambda g\dot{\phi}}{f}\frac{\psi^{3}}{3}+g^{2}\frac{\psi^{4}}{2}.$ (3.22)